Determine the mass number for atoms with the following subatomic makeups 10. 2 protons, 2 neutrons, and 2 electrons
1
Expert's answer
2021-10-05T23:58:00-0400

Both have approximately the same masses.


Absolute mass of neutron=1.675×10^(-27) ×2 kg


Relative mass of neutron=1 u


Absolute mass of proton=1.673×10^(-27) ×2kg


Relative mass of proton=1 u


Note: Since the mass of an electron is so less i.e. it is so negligible that it doesn't much affect the mass.


Mass of electron= 9.109×10^(-31) kg×2.